# Sri Lanka Document Datasets: A Large-Scale, Multilingual Resource for Law, News, and Policy (v20251005)

### v1

@@ Abstract
We present a collection of open, machine-readable document datasets covering parliamentary proceedings, legal judgments, government publications, news, and tourism statistics from Sri Lanka. As of v20251005, the collection currently comprises 215,670 documents (60.3 GB) across 13 datasets in Sinhala, Tamil, and English. The datasets are updated daily and mirrored on GitHub and Hugging Face. These resources aim to support research in computational linguistics, legal analytics, socio-political studies, and multilingual natural language processing. We describe the data sources, collection pipeline, formats, and potential use cases, while discussing licensing and ethical considerations.
Sri Lanka Document Datasets: A Large-Scale, Multilingual Resource for Law, News, and Policy (v20251005)
Nuwan I. Senaratna Independent Researcher nuwans@alumni.stanford.edu
@@ 3 Datasets
As of v20251005, Sri Lanka Document Datasets consists of 13 datasets.
@@
    Appeal Court Judgements: A Court of Appeal judgment is a higher court ruling that reviews decisions of lower courts, shaping legal precedent and protecting citizens’ rights. 10,146 documents, 10.3 GB, from 2012-04-23 to 2025-10-03. Source: https://courtofappeal.lk. Dataset: lk_appeal_court_judgements.
@@
    Supreme Court Judgements: A Supreme Court judgment is a binding legal decision that interprets the Constitution and laws, shaping justice, governance, and citizens’ rights. 1,575 documents, 1.3 GB, from 2016-07-22 to 2025-10-03. Source: https://supremecourt.lk. Dataset: lk_supreme_court_judgements.
@@
    Police Press Releases: A police press release is an official update from law enforcement on crimes, arrests, safety alerts, or public notices, ensuring transparency and public awareness. 736 documents, 250.4 MB, from 2025-05-01 to 2025-10-03. Source: https://www.police.lk. Dataset: lk_police_press_releases.
@@
    Acts: A legal act is a law passed by Parliament that governs rights, duties, economy, and society, shaping daily life and national policy. 3,928 documents, 6.8 GB, from 1981-01-22 to 2025-09-22. Source: https://documents.gov.lk. Dataset: lk_acts.
@@
    Extraordinary Gazettes: An Extraordinary Gazette is an official government publication used to announce urgent laws, regulations, or public notices with immediate effect. 101,532 documents, 20.1 GB, from 2010-01-01 to 2025-10-03. Source: https://documents.gov.lk. Dataset: lk_extraordinary_gazettes.
@@
    Cabinet Decisions: A Sri Lanka Cabinet Decision is an official policy or action agreed by the Cabinet of Ministers, shaping governance, law, and national development in the country. 10,369 documents, 125.3 MB, from 2010-09-27 to 2025-09-22. Source: https://www.cabinetoffice.gov.lk. Dataset: lk_cabinet_decisions.
@@
    Treasury Press Releases: A Sri Lanka Treasury press release shares key govt financial updates—on budgets, debt, or policy—vital for transparency, guiding investors, citizens, and markets on the nation’s economic direction. 133 documents, 142.9 MB, from 2015-09-08 to 2025-07-30. Source: https://www.treasury.gov.lk. Dataset: lk_treasury_press_releases.
@@
  11. 11.
    News: A collection of news documents. 79,167 documents, 1.2 GB, from 2021-09-12 to 2025-10-05. Source: multiple sources. Dataset: lk_news.
@@
    Tourism Weekly Reports: Report on Weekly Tourist Arrivals to Sri Lanka. 33 documents, 91.5 MB, from 2023-01-01 to 2025-09-01. Source: https://www.sltda.gov.lk. Dataset: lk_tourism_weekly_reports.

### v2

@@ Abstract
We present a collection of open, machine-readable document datasets covering parliamentary proceedings, legal judgments, government publications, news, and tourism statistics from Sri Lanka. The collection currently comprises of 229,858 documents (57.1 GB) across 24 datasets in Sinhala, Tamil, and English. The datasets are updated daily and mirrored on GitHub and Hugging Face. These resources aim to support research in computational linguistics, legal analytics, socio-political studies, and multilingual natural language processing. We describe the data sources, collection pipeline, formats, and potential use cases, while discussing licensing and ethical considerations. This manuscript is at version v2025-10-15-1111.
Sri Lanka Document Datasets: A Large-Scale, Multilingual Resource for Law, News, and Policy
Nuwan I. Senaratna Independent Researcher https://github.com/nuuuwan
@@ 3 Datasets
As of v2025-10-15-1111, Sri Lanka Document Datasets consists of 24 datasets and are publicly accessible on GitHub.  [^8] https://github.com/nuuuwan/lk_datasets
@@
    Appeal Court Judgements: A Court of Appeal judgment is a higher court ruling that reviews decisions of lower courts, shaping legal precedent and protecting citizens’ rights. 10,163 documents, 10.4 GB, from 2012-04-23 to 2025-10-14. Source: https://courtofappeal.lk. Dataset: lk_appeal_court_judgements.
@@
    Supreme Court Judgements: A Supreme Court judgment is a binding legal decision that interprets the Constitution and laws, shaping justice, governance, and citizens’ rights. 2,167 documents, 1.3 GB, from 2009-01-27 to 2025-10-13. Source: https://supremecourt.lk. Dataset: lk_supreme_court_judgements.
@@
    Police Press Releases: A police press release is an official update from law enforcement on crimes, arrests, safety alerts, or public notices, ensuring transparency and public awareness. 763 documents, 261.5 MB, from 2025-05-01 to 2025-10-14. Source: https://www.police.lk. Dataset: lk_police_press_releases.
@@
    Acts: A legal act is a law passed by Parliament that governs rights, duties, economy, and society, shaping daily life and national policy. 3,934 documents, 6.8 GB, from 1981-01-22 to 2025-10-07. Source: https://documents.gov.lk. Dataset: lk_acts.
@@
    Extraordinary Gazettes 2020S: An Extraordinary Gazette is an official government publication used to announce urgent laws, regulations, or public notices with immediate effect. 45,334 documents, 1.3 GB, from 2020-01-01 to 2025-10-11. Source: https://documents.gov.lk. Dataset: lk_extraordinary_gazettes_2020s.
@@
    Extraordinary Gazettes 2010S: An Extraordinary Gazette is an official government publication used to announce urgent laws, regulations, or public notices with immediate effect. 56,379 documents, 3.2 GB, from 2010-01-01 to 2019-12-31. Source: https://documents.gov.lk. Dataset: lk_extraordinary_gazettes_2010s.
@@
    Cabinet Decisions: A Sri Lanka Cabinet Decision is an official policy or action agreed by the Cabinet of Ministers, shaping governance, law, and national development in the country. 10,385 documents, 136.4 MB, from 2010-09-27 to 2025-10-07. Source: https://www.cabinetoffice.gov.lk. Dataset: lk_cabinet_decisions.
@@
    Treasury Press Releases: A Sri Lanka Treasury press release shares key govt financial updates—on budgets, debt, or policy—vital for transparency, guiding investors, citizens, and markets on the nation’s economic direction. 134 documents, 143.9 MB, from 2015-09-08 to 2025-10-07. Source: https://www.treasury.gov.lk. Dataset: lk_treasury_press_releases.
  11. 11.
@@
    News: A collection of news documents. 80,973 documents, 1.2 GB, from 2021-09-12 to 2025-10-15. Source: multiple sources. Dataset: lk_news.
@@
    Tourism Weekly Reports: Report on Weekly Tourist Arrivals to Sri Lanka. 34 documents, 96.5 MB, from 2023-01-01 to 2025-10-01. Source: https://www.sltda.gov.lk. Dataset: lk_tourism_weekly_reports.
  14. 14.
@@
  15. 15.
    Dmc Situation Reports: Situation Report including information about Heavy Rain, Wind, Tree Falling, Lighting etc. 4,321 documents, 3.0 GB, from 2018-01-02 to 2025-10-15. Source: https://www.dmc.gov.lk. Dataset: lk_dmc_situation_reports.
  16. 16.
    Dmc Weather Forecasts: Weather Forecasts for various places in Sri Lanka. 3,655 documents, 4.3 GB, from 2023-03-26 to 2025-10-15. Source: https://www.dmc.gov.lk. Dataset: lk_dmc_weather_forecasts.
  17. 17.
    Dmc River Water Level And Flood Warnings: River Water Level and Flood Warnings for various places in Sri Lanka. 21 documents, 7.1 MB, from 2025-06-10 to 2025-10-15. Source: https://www.dmc.gov.lk. Dataset: lk_dmc_river_water_level_and_flood_warnings.
  18. 18.
    Dmc Landslide Warnings: Landslide Warnings including early warnings, locations of potential risk, areas and places which need special attention, and automated landslide early warning map. 564 documents, 436.7 MB, from 2019-09-26 to 2025-10-13. Source: https://www.dmc.gov.lk. Dataset: lk_dmc_landslide_warnings.
  19. 19.
    Cbsl Annual Reports: Annual Reports of the Central Bank of Sri Lanka (CBSL).It has been discountinued since 2023and replaced with swo separate reports,namely, the Annual Economic Reviewand Financial Statementsand Operations of the Central Bank. 1,137 documents, 3.5 GB, from 1950-01-01 to 2023-01-01. Source: https://www.cbsl.gov.lk. Dataset: cbsl_annual_reports.
  20. 20.
    Fisheries Annual Statistics Reports: Annual Fisheries Statistics Reports of the Ministry of Fisheries,Aquatic and Ocean Resources, Sri Lanka 9 documents, 15.1 MB, from 2017-01-01 to 2024-01-01. Source: https://www.fisheries.gov.lk. Dataset: lk_fisheries_annual_statistics_reports.
  21. 21.
    Fisheries Monthly Fish Production Reports: Monthly Fish Production Reports of the Ministry of Fisheries,Aquatic and Ocean Resources, Sri Lanka 96 documents, 8.9 MB, from 2019-01-01 to 2025-08-01. Source: https://www.fisheries.gov.lk. Dataset: lk_fisheries_monthly_fish_production_reports.
  22. 22.
    Fisheries Weekly Fish Prices Reports: Weekly Fish Prices Reports of the Ministry of Fisheries,Aquatic and Ocean Resources, Sri Lanka 248 documents, 21.8 MB, from 2019-01-01 to 2025-09-22. Source: https://www.fisheries.gov.lk. Dataset: lk_fisheries_weekly_fish_prices_reports.
  23. 23.
    Fisheries Monthly Export Import Reports: Monthly Fish Export and Import Reports of the Ministry of Fisheries,Aquatic and Ocean Resources, Sri Lanka 64 documents, 51.6 MB, from 2019-01-01 to 2025-08-01. Source: https://www.fisheries.gov.lk. Dataset: lk_fisheries_monthly_export_import_reports.
  24. 24.
    Edupub: Educational Publications from the Educational Publications Department, Sri Lanka. 1,426 documents, 866.2 MB, from 2025-01-01 to 2025-01-01. Source: http://www.edupub.gov.lk. Dataset: lk_edupub.